OpenStudy (anonymous):

Can someone walk me through how to switch the order of integration on some double integrals?

OpenStudy (anonymous):

\[\text{Original integral:}~\int_0^1\int_{y/2}^{1/2}e^{-x^2}dxdy\]

OpenStudy (anonymous):

I tried changing the bounds, and I got \[\int_0^{1/2}\int_0^{2x}f(x,y)dydx\]Am I on the right track so far?

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (anonymous):

Im kind of confused on what to do next though.

ganeshie8 (ganeshie8):

next integrate

ganeshie8 (ganeshie8):

start with the inner integral : \[\large \int_0^{1/2} \color{blue}{\int_0^{2x}e^{-x^2}dy}~dx\]

OpenStudy (anonymous):

OH! I forgot that I can actually integrate it now. So, it becomes: \[\int_0^{1/2}2xe^{-x^2}dx=-(e^{-x^2})|_0^{1/2}=1-e^{-1/4}\]Is that right?

ganeshie8 (ganeshie8):

Perfect! you may use wolfram to double check http://www.wolframalpha.com/input/?i=%5Cint_0%5E1%5Cint_%7By%2F2%7D%5E%7B1%2F2%7De%5E%7B-x%5E2%7Ddxdy

OpenStudy (anonymous):

Alright, thanks! I thought I had to do something to the integrand, and I didn't recognize that with the change to dydx, I could actually integrate.

ganeshie8 (ganeshie8):

Thats it! changing order of integration gave us that factor 2x which was useful in u-substitution

OpenStudy (anonymous):

Thanks. :)

ganeshie8 (ganeshie8):

np:)